OpenStudy (anonymous):

With the $5,000 he has to spend, Jim buys 5,000 carnations at $0.30 each, 4,000 tulips at $0.60 each, and 300 irises at $0.25 each. Write an inequality which describes how many roses, r, Jimcan buy if roses cost $0.80 each.

OpenStudy (texaschic101):

5000(.30) + 4000(.60) + 300(.25) + r(.80) <= 5000 1500 + 2400 + 75 + .80r <= 5000 3975 + .80r <= 5000 .80r <= 5000 - 3975 .80r <= 1025 r <= 1025/.80 r <= 1281.25